OpenStudy (anonymous):

Given the plane that passes through the points (3, -4, -6), (5, -9, 10) and (-5, -7, -2). Let A(x0,y0,z0) be a point on the plane with x0 = 0 and y0 = 3. Find z0

OpenStudy (amistre64):

this we can cross to get a normal

OpenStudy (amistre64):

lets set a point at origin to determine the vectors involved: (3, -4, -6) (5, -9, 10) (-5, -7, -2) -3 +4 +6 -3 +4 +6 -3 +4 +6 --------------------------------- (0, 0, 0) <2,-5, 16> <-8,-3,4>

OpenStudy (amistre64):

do you see why id did that?

OpenStudy (amistre64):

when we can measure the points from the origin, we can turn them into vectors

OpenStudy (anonymous):

yeah just titing the triangle so that 1 side starts at the origin. what does that do tho

OpenStudy (amistre64):

< 2,-5, 16> <-8,-3, 4 > ----------- x = -20+48 = 28 y = -128-8 = -136 z = -6-40 = -46 check the math on that

OpenStudy (anonymous):

you can turn 2 of them into vectors well erm

OpenStudy (anonymous):

? where are those numbers coming from

OpenStudy (amistre64):

im crossing the 2 vectors i found from the points

OpenStudy (amistre64):



OpenStudy (amistre64):

a x b = normal to the plane

OpenStudy (amistre64):

tell me where any of this doesnt make sense

OpenStudy (anonymous):

nah i understand what you are doing graphically but where does this come from x = -20+48 = 28 y = -128-8 = -136 z = -6-40 = -46

OpenStudy (amistre64):

i found my 2 vectors that reside in the plane: < 2,-5, 16> <-8,-3, 4 > ; given 2 vectors in the plane, if you cross them you get the normal that is perp to the plane

OpenStudy (amistre64):